OpenStudy (anonymous):

Which polynomial is a special product and why? x²-81or x²-13x+36.

OpenStudy (anonymous):

ithink both because both of them we can write interms of x²-81=x²-9^2 =(x+9)(x-9) x²-13x+36.= x²-9x-4x+36.=x(x-9)-4(x-9) =(x-4)(x-9)